OpenStudy (anonymous):

6. A particular model of walkie-talkie can broadcast in a circular area. The radius of the broadcast area is 10,000 feet. Find the area of this circle to the nearest square foot. Use 3.14 for pi.

OpenStudy (apoorvk):

pi r^2 = area = (10000)^2 x 3.14 = 3.14 x 10^8

OpenStudy (anonymous):

For this, you would want the formula for the area of a circle: \[A = \pi r ^2 \] If the radius of the broadcast area is 10,000 feet, then you can use the formula and replace r with 10,000. \[A = \pi (10,000) ^2 \] They specifically asked you to use "3.14" for pi. \[A = (3.14) (10,000) ^2 \] Order of operations would indicate that you need to deal with the power of 2 on the 10,000 first. \[A = (3.14) (100,000,000)\] From here, you can move the decimal in the 3.14 to the right 8 places (one place for each zero in the 100,000,000). \[A = 314,000,000\] When you answer the question with a complete sentence, don't forget to put the units. The area of the broadcast is \[314,000,000 ft ^{2}\]